### v1

# Bruno: Backpropagation Running Undersampled for Novel device Optimization

Abstract
Recent efforts to improve the efficiency of neuromorphic and machine learning systems have focused on the development of application-specific integrated circuits, which provides hardware specialized for the deployment of neural networks, leading to potential gains in efficiency and performance. These systems typically feature an architecture that goes beyond the von Neumann architecture employed in general-purpose hardware such as graphic processing units. Neural networks developed for this specialised hardware then needs to take into account the specifics of the hardware platform, which requires novel training algorithms and accurate models of the hardware, since they cannot be abstracted as a general-purpose computing platform. In this work, we present a bottom-up approach to train neural networks for hardware based on spiking neurons and synapses built on ferroelectric capacitor (FeCap) and Resistive switching non-volatile devices (RRAM) respectively. In contrast to the more common approach of designing hardware to fit existing abstract neuron or synapse models, this approach starts with compact models of the physical device to model the computational primitive of the neurons. Based on these models, a training algorithm is developed that can reliably backpropagate through these physical models, even when applying common hardware limitations, such as stochasticity, variability, and low bit precision. The training algorithm is then tested on a spatio-temporal dataset with a network composed of quantized synapses based on RRAM and ferroelectric leaky integrate-and-fire (FeLIF) neurons. The performance of the network is compared with different networks composed of leaky integrate-and-fire (LIF) neurons. The results of the experiments show the potential advantage of using BRUNO to train networks with FeLIF neurons, by achieving a reduction in both time and memory for detecting spatio-temporal patterns with quantized synapses.
Keywords Learning  $\cdot$ FeCAP  $\cdot$ neuromorphic  $\cdot$ SNN  $\cdot$ ReRAM
1 Introduction
The rapid growth of artificial intelligence (AI) demands efficient hardware solutions that enable energy-efficient computing without sacrificing performance. However, current AI hardware faces fundamental trade-offs between power consumption, computational efficiency, and accuracy. Furthermore, the storage and routing of synaptic weights remains one of the major challenges for efficient hardware platforms for Spiking Neural Networks [1]. While advantageous, quantization of synaptic weights can introduce noise and nonlinearity that degrade the network’s accuracy performance in inference. Neurons using innovative devices promise better noise rejection [2] and, as we will show, allow strict weight quantization. Moreover, memory devices have shown potential in their application in Machine Learning (ML) and neuromorphic engineering for in-memory computing [3] due to their non-volatility, compact size, and low latency [4]. Thanks to these properties, devices such as RRAM, phase-change memory (PCM) or ferroelectric devices offer situational advantages over traditional memories such as static random-access memory (SRAM). Their non-volatility makes it possible to store the parameters of ML models without a constant power supply. This characteristic makes them ideal for representing static and plastic synapses of an artificial neural network.
However, the use of these memory devices within an ASICs requires accurate physical or behavioural models. These models typically represent the network synapses, where their behavioural models are incorporated into simulation frameworks and training algorithms [5, 6, 7, 8, 9, 10] to account for device variability in the updating process. Recent research has also incorporated physical models into machine learning frameworks such as AIHWKIT [4, 11]. However, these devices can also be used as computational primitives in the neuron to enhance its behaviour by including multiple time constants [2]. When used as computational primitives, these models have to be computationally efficient and well-posed to allow efficient temporal training methods, such as backpropagation through time (BPTT) [12]. Accurate models and efficient training are particularly important when the inherent physical dynamics of the underlying hardware are to be exploited. Algorithms and hardware need to be co-developed from memory devices and circuit technology to the whole neuromorphic system.
One of the challenges of using memory devices is that they are inherently stateful and exhibit non-linear behaviour over a wide range of timescales that need to be reliably modelled. Given that typical application timescales are on the order of 100 milliseconds to seconds [13, 14, 15], integrating these models, whose simulation timescale is on the order of microseconds, with gradient-based training algorithms, such as BPTT, would lead to several problematic scenarios. Firstly, there could be vanishing and exploding gradient problems due to the large number of simulation time steps. Secondly, the high number of time steps significantly increases the training time and memory requirements, as the computational graph created for the automatic differentiation contains the states at every microsecond. An alternative approach would be to increase the time step, but this would lead to model instability and inaccurate results. Therefore, we propose an innovative training method that reduces the training time by around an order of magnitude and reduces the time and memory consumption during training and inference over a large number of time steps. This is achieved by using two different time resolutions for the forward and the backward passes. We call this method Backpropagation Running Undersampled for Novel device Optimization (BRUNO). This training method has several advantages, allowing us to train neural networks with high temporal dynamics, and compare the behaviour of networks with these new models with standard neurons such as LIF [16]. To validate BRUNO, a network composed of FeLIF neurons and synapses based on RRAM devices is trained and compared with LIF neurons. In addition, we show how the FeLIF model could outperform or match recurrent neural networks when the synaptic weights are of low bit precision.
2 Materials and Methods
2.1 Ferroelectric LIF neuron
The FeLIF neuron [2], which co-integrates complementary metal-oxide semiconductor (CMOS) and FeCap technology, uses the non-linear charge-voltage relationship of the FeCap to integrate the inputs to two different state variables, the dielectric and the ferroelectric polarization. The dielectric polarization is linear with the voltage across the capacitor and would result in the behaviour of a conventional LIF neuron. However, the non-volatile ferroelectric polarization is gated by the membrane potential Vmem, the voltage across the capacitor, and only increases when a certain membrane potential Vc is reached. The ferroelectric polarization then accumulates until the saturation polarization Ps is reached. The FeLIF can thus be described phenomenologically by an integrate-and-fire (IF) neuron gated by a leaky integrator (LI) (Figure 1(a)), resulting in a neuron capable of tracking significant inputs over long time frames.
Figure 1: (a) Operating principle of the FeLIF neuron. When the membrane potential V is charged by the input current Isyn via the leaky integrator (LI) and reaches the coercive voltage Vc, the current partially flows to the integrator (I). When the integrator state P reaches the saturation value Ps, the current flows back to the leaky integrator until it reaches the firing threshold and the neuron emits a spike. (b) Polarization and voltage of the neuron for a constant input current. As described in the operating principle, the input current is integrated into the polarization as long as $V\mathrm{{}_{mem}}>V\mathrm{{}_{c}}$ and $P<P\mathrm{{}_{s}}$
Based on the model extracted from the FeCap-CMOS neuron circuit [2], the behaviour can be described as follows. The derivative of the membrane potential is obtained by dividing all the currents flowing into the FeCap by its capacitance and the parasitic capacitance originating from the CMOS circuit:
$\frac{dV\mathrm{{}_{mem}}}{dt}=\frac{I_{\mathrm{syn}}-I_{\mathrm{leak}}-I_{p}}{C_{\mathrm{0}}+C_{\mathrm{par}}}$ (1)
similar to a LIF neuron, where Vmem is the membrane potential, $\mathrm{I_{syn}}$ is the synaptic input current, $\mathrm{I_{leak}}$ is the leakage current through the ferroelectric capacitor and $\mathrm{I_{p}}$ is the displacement current of the ferroelectric polarization. The ferroelectric displacement current results from the ferroelectric polarization process under an electric field and is proportional to the derivative of the polarization:
$\frac{I_{p}}{A}=\frac{dP}{dt}=\frac{\textrm{sign}(E_{fe})\cdot P_{\textrm{s}}-P}{\tau(E_{fe})}$ (2)
where $A$ is the area of the device, $\mathrm{P}$ is the ferroelectric polarization, $E_{fe}$ is the electric field across the ferroelectric layer and $\mathrm{P_{s}}$ is the saturation polarization. $\mathrm{\tau}$ is the time constant of the polarization and is a function of the electric field:
$\tau(E)=\tau_{0}\cdot\exp\left(\frac{E_{\textrm{a}}}{|E_{fe}|}\right)^{\alpha}$ (3)
with activation electric field $\mathrm{E_{a}}$, elementary time constant $\mathrm{\tau_{0}}$ and the fitting factor $\mathrm{\alpha}$. The parameters described above used for the simulations are listed in Table 1.
Parameter Value
A 25 µm2
C0 0.558 pF
Cpar 15 fF
PS 0.22 C/m2
Ea 1.27 V/nm
$\tau_{0}$ 0.1 ps
$\alpha$ 1.3
Table 1: Physical parameters of the FeLIF neuron model used for simulation.
2.2 Resistive switching non-volatile synapses
In this work, the weight quantization is motivated by the use of RRAM as n-bit synaptic connections within the network. The RRAM switching behaviour is simulated using the Jülich Aachen Resistive Switching Tool (JART) valence change mechanism (VCM) physical compact model [17]. A detailed description of the model and the physical stack of the simulated RRAM can be found in the Supplementary Information, Section 1.1.
In this work, current-pulse programming, a common approach in neuromorphic hardware [18, 19], is used to program the RRAM device. Therefore, the analogue behaviour of the RRAM is studied with current pulses of 10 µs time width and amplitudes from 50 µA to 300 µA with 10 µA steps. After each programming pulse, a read-out operation is performed in current mode. During this read-out operation, a current pulse with 10 µA amplitude and 10 µs pulse width is applied to the RRAM and the voltage drop across the cell is measured to calculate the corresponding conductance.
To qualitatively discuss the impact of RRAM stochasticity on device programming, the filament-related parameters responsible for the conductance variation are varied by 1 $\mathrm{\sigma}$ from their mean values, as shown in the Supplementary Information, Table 1.
The quantized values of the RRAM were obtained with current pulses of 10 µs time width and amplitudes as in Table 2. Note that these values were obtained with partial set operations starting from the high resistive state (HRS) and reaching different intermediate low resistive states (LRS).
States Current pulse amplitude
$HRS$ 10 µA
$LRS_{1}$ 50 µA
$LRS_{2}$ 80 µA
$LRS_{3}$ 110 µA
$LRS_{4}$ 140 µA
$LRS_{5}$ 180 µA
$LRS_{6}$ 230 µA
$LRS_{7}$ 300 µA
Table 2: Current pulse amplitudes for 3-bits quantization
2.3 Quantization-Aware training
The quantization of the synapses throughout the training process has to be considered to adequately model the RRAM-based synapses. Weight quantization is incorporated into the training phase using the Quantization Aware Training (QAT) process [20]. Since the quantization operation is non-differentiable, a surrogate quantization operation is used to perform the weight quantization. This means that quantization is applied in the forward pass, but is replaced by an approximate differentiable function in the backward pass.
$w_{q}=\mathrm{sround}\left(\frac{w}{s_{w}}\right)$ (4)
$s_{w}=\frac{\mathrm{max}\left(|w|\right)}{2^{N-1}-1}$ (5)
The quantization operator is defined in the equation 4, where the trained floating-point weight value $w$ is scaled by a factor $s_{w}$ defined in the equation 5, where N is the number of bits to be considered. The scaled weight is then applied to a stochastic rounding operator [21]. Since the gradient of the rounding is approximately zero over the domain, the straight-through estimators (STE) surrogate gradient is applied to it [22, 23].
3 Results
3.1 Training algorithm
Training physically accurate neuron models with different temporal dynamics can cause instabilities at simulation timescales of $\approx$1 ms. In addition, training at finer timescales of around 1 µs can cause gradient optimization problems and increase computational resources by three orders of magnitude. BRUNO uses a dual timescale for the forward pass and the gradient computation. Figure 2 illustrates the training procedure. In the forward pass, the network simulation runs with a fine-grained simulation time step of 1 µs. However, the gradient calculation is performed with a 1 ms time step. This reduces the size of the unrolled computational graph during backpropagation by reducing the total number of time steps in the backward pass. In addition, the values used to compute the gradient are obtained during the 1 µs time step duration, making the training more stable than training with 1 ms in the forward pass. To achieve the dual timescale training, the calculation of the internal state of the neuron can be computed on both time steps and use the higher timescale gradient as a surrogate gradient for the 1 µs time resolution value.
Figure 2: Schematic of the algorithm for computing the forward and backward passes with different time steps. During the forward pass (blue box) the simulation operates on 1 µs timescale. In the backward pass (green box) the gradient is calculated on 1 ms timescale.
To validate BRUNO, the FeLIF model described in section 2.1 is used. The internal state of the neuron $s^{t}=\{V^{t},P^{t}\}$ is defined as the voltage and polarization over time. Algorithm 1 shows the pseudocode of the training method. In the forward pass, the value of the neuron state at time t ($s^{t}$) is calculated with microsecond ($s_{\mu s}$) and millisecond ($s_{ms}$) time resolution. By applying the function $s_{ms}+detach(s_{\mu s}-s_{ms})$, the state of the neuron is equal to $s_{\mu s}$. However, as the subtraction $s_{\mu s}-s_{ms}$ is detached from the computational graph, the gradient flows throughout the millisecond time scale state ($s_{ms}$). The final voltage $v_{mem}$ obtained from $s_{ms}$ is then applied to a threshold operator $sg$ that applies a surrogate gradient [24]. When the neuron fires, the state is reset.
Algorithm 1 BRUNO
for $t=1,\dots,T$ do
     $s_{\mu s}^{0}\leftarrow s^{t-1}$
     for $i=1,\dots,1000$ do
         $s_{\mu s}^{i}\leftarrow$ compute $s$ based on equations 1 - 3 with 1 µs $\Delta$t
     end for
     $s_{ms}\leftarrow$ compute $s$ based on equations 1 - 3 with 1 ms $\Delta$t
     $s^{t}\leftarrow s_{ms}+detach(s_{\mu s}-s_{ms})$
     spike $\leftarrow sg(v_{mem}^{t}-v_{thr})$
     reset $s^{t}$ in case of spike
end for
3.2 Verification and benchmarking
To verify the functionality of BRUNO on a physically accurate neuron model for future hardware deployment, first, the behaviour of the FeLIF Python model used during training is verified to match the SPICE simulations of the FeLIF circuit. Second, the RRAM stochasticity and quantization for the synapse implementation are analysed. Finally, a SNN is trained on a spatio-temporal benchmark using the obtained quantization values, and the results are compared between different cases, namely training with and without BRUNO, using different quantization levels for the weights and employing different neuron models.
3.2.1 Hardware model and verification
A comparison of the SPICE circuit simulations and the behavioural model (equations 1-3) shows a close match in the transient of the ferroelectric polarization (Figure 3(a)) as well as in the membrane potential of the neuron (Figure 3(b)). While the network simulation resets the membrane potential as soon as the firing threshold is crossed, the circuit needs to reset the physical device, resulting in an overshoot in the voltage. However, since it happens during the refractory period, it does not affect the operation of the neuron in the network simulations.
Figure 3: Comparison of the membrane potential transient between the SPICE simulation of the neuron circuit and the model used in the network simulations with a DC input current of 308 pA. Except for the spike amplitude, the polarization (a) and the voltage (b) match between the Python and SPICE simulations. The amplitude of the spike in the network is negligible as it is registered as an event in any case.
3.2.2 Synaptic Stochastic and Quantization analysis
Due to their non-volatility and small footprint, RRAMs were chosen as the artificial synapses for the neural networks used in the BRUNO benchmarking. They operate on the principle of the formation or rupture of conductive filaments, resulting in the switching between LRS and HRS. However, due to the physical processes leading to this switching (the random formation and rupture of the conductive filaments during the SET and RESET processes [25]), RRAM devices exhibit inherent stochasticity, leading to Cycle-to-Cycle (C2C) variability. As a result, adjacent conductance states may overlap, making them indistinguishable. Therefore, quantization is required to define a limited number of discrete, non-overlapping conductance levels to ensure distinguishable weights/conductance states. The inherent stochastic behaviour of RRAM devices was incorporated into the device simulations using stochastic rounding as described in Section 2.3.
Before quantizing the RRAM programmed states, it is essential to investigate the RRAM analogue weight update behaviour and assess its linearity. Figure 4(a) illustrates the change in conductance in response to an increase in the amplitude of the programmed current pulse, as well as the maximum variability of the programmed conductance state due to the stochastic switching behaviour [26, 27]. It can be observed that a quasi-linear change in conductance follows the application of each programming pulse, with an increase in the amplitude of the current pulse. However, the C2C variability limits the number of states that can be reliably read, since the same analogue conductance state can be obtained with different current values. As a result, the distributions of different states overlap, preventing a correct reading. To reliably extract distinguishable RRAM programmed states from the quasi-analogue weight update, state quantization is performed. Based on the data in Figure 4(a), 3-bit states have been quantized. Figure 4(b) illustrates the distribution of 3-bit quantized states obtained from Monte Carlo (MC) simulations, which included 7 LRS states extracted from the analogue weight update curve and 1 initial HRS state, as explained in Section 2.2. The distribution of conductance within 2 $\mathrm{\sigma}$ over the selected programming pulses shows a clear distribution across the entire conductance range. The distribution of adjacent conductance states shows a distinct gap with an average of 20 µS, which is sufficient to be detected by standard CMOS circuitry.
Figure 4: RRAM state programming and device conductance. (a) Device conductance based on the programming current applied, for a mean filament size of 0.4 nm length and 45 nm filament radius, an upper limit of 0.36 nm length and 49 nm filament radius, and a lower limit of 0.44 nm length and 41 nm filament radius. (b) Distribution of the different 3-bit quantized states.
3.2.3 Benchmark
To test the dual time constant of the FeLIF neuron for temporal information processing, spatio-temporal data from the Neuromorphic Braille Reading Dataset [28] is used. The dataset consists of 5,400 braille reading samples recorded by an Omega 3 robot equipped with a sensorized fingertip over 3D-printed braille letters [28].
Figure 5: Network architectures used in the benchmarks. All networks contain an encoding layer, a hidden layer with 256 LIF neurons, and an output classification layer. (a) Feedforward network with FeLIF neuron in the output layer. (b) Feedforward network with the LIF neurons in the output layer. (c) Network where the LIF hidden layer has an explicit recurrent connection and the output neurons are LIF neurons.
To evaluate the performance of the FeLIF neuron and BRUNO, three different networks were constructed. The three networks have an input encoding layer, which transforms the analogue sensor values of the Braille letters into spikes. They also include a hidden layer of 256 LIF neurons and an output layer of 27 neurons, representing each of the Braille letters, including the space. The first network uses FeLIF neurons in the output layers (Figure 5(a)), while the second uses LIF neurons (Figure 5(b)). Moreover, for spatio-temporal data, LIF neuron networks may have difficulty learning long temporal dependencies due to the short-term dynamics [29]. Adding explicit recurrence can improve the performance on spatio-temporal data [30]. For this reason, the third network incorporates recurrent connection weights in the hidden layer (Figure 5(c)). A Hyperparameters Optimization (HPO) was performed to train the three network architectures SNN with different synaptic quantization values for the RRAM. The final parameters used can be found in Table 3.
To BRUNO or not to BRUNO.
The first comparison made was between training a SNN with and without BRUNO. Due to the reduction of the computational graph in the backward pass, the training time of the SNN is also reduced. The tests were performed with the parameters in Table 3 and floating-point precision. As a result, Figure 6 shows the time for training and inference on CPU of a FeLIF network with 64, 128, and 256 neurons, for sequences from 10 to 500 ms. The training of the network for sequences of 1000 ms with BRUNO reduces the amount of memory requirements from 5.3 GB to 15.3 MB, 0.28% of the memory consumption of the vanilla method. In addition, the vanilla training method led to the occurrence of exploding gradients during training. In contrast, BRUNO demonstrated the ability to train the neural network with a final accuracy of $91.34\%\pm 0.59$ for full-precision weights.
Figure 6: Time analysis between using or not BRUNO during inference and training.
Figure 7: Memory analysis between using or not BRUNO during inference and training.
Quantization and model comparison
The next validation performed was the comparison between different synaptic quantization levels and neuron models. As explained earlier, three different networks were created comparing the FeLIF neuron model with a feedforward and recurrent LIF. For each network and quantization value, HPO was performed using the Optuna library [31]. Once the optimal hyperparameters were obtained (Table 3) for each experiment (network and quantization), 50 simulations were performed with different initialisation values. Figure 8 and Table 4 show the final accuracy obtained. Without quantization, the feedforward FeLIF network achieves a similar performance as the feedforward LIF network with an accuracy of $91.44\%\pm 0.73$ compared to $91.34\%\pm 0.59$ for the FeLIF. In contrast, the recurrent network has an accuracy of $95.30\%\pm 0.61$. However, as the quantization level of the network increases, the performance of the feedforward FeLIF becomes better than that of the LIF with recurrence, obtaining $74.73\%\pm 1.37$ and $74.26\%\pm 2.48$ respectively, compared to the feedforward LIF which drops drastically in performance to $40.48\%\pm 27.84$.
Quant. Model $\alpha_{hid}$ $\beta_{hid}$ $\alpha_{out}$ $\beta_{out}$ LR [^1] Learning rate FeLIF Vthr
3 LIF 0.662 0.703 0.565 0.696 3.429E-02 -
RLIF 0.603 0.310 0.354 0.295 1.478E-03 -
FeLIF 0.468 0.735 - - 7.840E-04 3.039
4 LIF 0.226 0.245 0.865 0.834 3.276E-03 -
RLIF 0.599 0.710 0.390 0.349 6.055E-03 -
FeLIF 0.456 0.322 - - 3.716E-03 2.544
8 LIF 0.230 0.302 0.901 0.591 1.283E-02 -
RLIF 0.186 0.537 0.347 0.846 6.766E-03 -
FeLIF 0.682 0.662 - - 4.387E-03 2.928
FP LIF 0.959 0.202 0.716 0.600 3.006E-03 -
RLIF 0.343 0.406 0.764 0.874 6.174E-03 -
FeLIF 0.299 0.147 - - 2.628E-03 3.388
Table 3: Parameters used for network simulations with different quantization levels and models. The $alpha_{hid/out}$ parameter is the membrane decay constant for the hidden and output layers. The $\beta_{hid/out}$ parameters are the synapse decay constants for the hidden and output layers. The FeLIF Vthr parameter is the firing threshold of the FeLIF neuron.
Network FP32 8 bits 4 bits 3 bits
$\mu$ $\sigma$ $\mu$ $\sigma$ $\mu$ $\sigma$ $\mu$ $\sigma$
LIF 91.44 0.73 91.81 0.73 78.10 1.85 40.48 27.84
RLIF 95.30 0.61 95.29 0.59 91.96 0.84 74.26 2.48
FeLIF 91.34 0.59 90.69 0.57 84.49 1.22 74.73 1.37
Table 4: Test accuracy obtained for different quantization levels and network architectures on the neuromorphic Braille letter reading dataset.
Figure 8: Accuracy Braille dataset with different quantization levels and networks. Each plot has different quantization levels (Floating-point, 8, 4, and 3 bits). In each plot, the three different network architectures are shown, trained with different initialisation seeds using the hyperparameters obtained using HPO
4 Discussion and conclusion
This work presents a method for training high-temporal neuron models that describe the physical accuracy behaviour of analogue hardware. The simulation of these models can exhibit instability when matching application time steps of $\approx$1 ms. The proposed training method allows the implementation of standard techniques, such as BPTT in the training of the FeLIF model, thereby reducing the training time by an order of magnitude and overcoming problems associated with exploding gradients. The FeLIF neuron model was used to validate the training algorithm. The model was trained with and without BRUNO on the Braille reading dataset, a dataset for benchmarking spatio-temporal patterns. The results show that training with BRUNO improves training time, memory requirements, and accuracy. Achieving a reduction of the 0.28% of the peak memory usage for training long temporal sequences ¿1 s. With an accuracy of $91.34\%\pm 0.59$ on the braille dataset. In contrast, training with vanilla BPTT produces gradient optimisation problems that make it impossible to train the network.
In addition to the validation of the training procedure with the vanilla method, benchmarking was performed with feedforward and recurrent LIF networks. This benchmarking was also performed with synaptic quantization for its possible future implementation on hardware with RRAM synapses. Different quantization levels were performed, for 8, 4, and 3 bits, where Hyperparameters Optimization (HPO) was performed to obtain the optimal set of hyperparameters for each network and quantization level. As a result of this benchmark, we show that a feedforward FeLIF network has comparable results to a feedforward LIF network with $91.34\%\pm 0.59$ and $91.44\%\pm 0.73$ respectively. However, for the quantization network, the FeLIF outperforms the feedforward LIF network and has similar performance with a recurrent LIF network with 3-bit quantization with $74.73\%\pm 1.37$ and $74.23\%\pm 2.48$ respectively. In contrast, the mean accuracy obtained for the feedforward network is $40.48\%\pm 27.84$, which could be due to a high learning rate for the quantized weights, since the network started to lose performance drastically on the last epochs.
BRUNO opens up the possibility of simulating and training neural models with complex dynamics that can be used for spatio-temporal pattern recognition. These neural models have the potential to be integrated into hardware with synaptic quantization, which could provide performance comparable to that of a recurrent LIF network, while requiring fewer resources and trainable parameters.
Acknowledgements
We thank Pedro F. Ortiz Costa and Dr. Francesco Malandrino for their support and fruitful discussions. We thank the Center for Information Technology of the University of Groningen for their support and for providing access to the Hábrók high performance computing cluster. This work was partially supported by the European Research Council (ERC) through the European’s Union Horizon Europe Research and Innovation Programme under Grant Agreement No 101042585. Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. The authors would like to acknowledge the financial support of the CogniGron research center and the Ubbo Emmius Funds (Univ. of Groningen).